OpenStudy (anonymous):

Medals will be given need help with problems, have the answers already just need to work them out. 7.3x^2 + 11x -4 9. 15x^2-55x-20 10. x^2-8x + 16 11. x^2 - 16 12.16^2-121 13. 125a^2 - 121 14. 9z^3 + 18z^2 -7z-14

OpenStudy (anonymous):

@surjithayer can you help please with these

OpenStudy (anonymous):

7. 3*-4=-12 11=12-1 12*-1=-12 9. 15*-20=-300 55=60-5 60*-5=-300 10. 1*16=16 8=4+4 4*4=16 11.12.13. (mx)^2-(ny)^2=(mx+ny)(mx-ny) 14. 9z^3+18z^2-7z-14=9z^2(z+2)-7(z+2) =(z+2)(9z^2-7) =(z+2){(3z)^2-(sqrt7)^2} =(z+2)(3z+sqrt7)(3z-sqrt7)

OpenStudy (anonymous):

omggggg thank you so much your the best dude, so when you fine the factor do you multiply ?

OpenStudy (anonymous):

this is the proper way to find the factos.

OpenStudy (anonymous):

okay thank you i understand you were a biggg help!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@surjithayer whats 11 12 13 ?

OpenStudy (anonymous):

nervermind i can do it myself lolll

OpenStudy (anonymous):

11,12,13 are questions. They are solved like i have solved in last part of 14.